the supplement of an angle is thirty degrees less than twice the measure of the angle itself.find the angle and its supplements.
Answer by BurningFlame_marc(77)

Hello!

First let's solve this step by step..

Representation:
let x= angle, 2x-30= the supplement.
Equation:
x+(2x-30)=180
Solution:
x+2x-30=180
3x=210
x=70

Final Answer:
the angle is 70 and the supplement is 110.